If a=(cos2Π)/7+(isin2Π)/7 .Then quadratic equation whose roots are α=a+a 2 +a 4 and β=a 3 +a 5 +a 6 is

If    a=(cos2Π)/7+(isin2Π)/7     .Then quadratic equation whose roots are  α=a+a2+a4 and β=a3+a5+a6 is





Grade:11

1 Answers

Ashwin Muralidharan IIT Madras
12 years ago

Hi Dushyant,

 

This is a bit tricky question. The idea behind the solution which I give you here can be used in many places to get the quadratic equation, whose roots satisfy some given conditions.

 

First of all: a^0 + a^1 +......+ a^6 = (a^7 - 1)/(a - 1) [using the sum of GP formula]

But a^7 = cos(2pi) + i*sin(2pi) = 1..... Hence the above sum = 0.

 

For the sake of typing, let me use "p" in place of alpha and "q" in place of beta.

 

So sum of roots = p+q = a + a^2 +.....a^6 = -a^0 = -1.

Next p^2 = (a + a^2 + a^4)^2 = a^2 + a^4 + a^8 + 2(a^3 + a^5 + a^6) ---------------- (1)

Also note that a^8 = a^7*a = a [since a^7 = 1]

 

Hence you have p^2 = p + 2(q) --------------[from (1)]

But we also have shown that p+q = -1, so q = -1-p

Hence p^2 = p -2*(1+p)

ie p^2 + p + 2 = 0......... (similarly you can also check up that q^2 + q + 2 = 0, by substitution of p = -1-q in the equation)

 

Hence both p & q satisfy the equation, x^2 + x + 2 = 0 (where p & q are the roots of that equation).....

And that solves the problem.

 

Hope that helps.

All the best. In case of any other questions feel free to mail me at [ashwin.rsmd@gmail.com]

 

Regards,

Ashwin (IIT Madras)